# There and Back Again: On the relation between Noise and Image Inversions in Diffusion Models

### v4

@@ 1 Introduction
Diffusion-based probabilistic models (DMs), (Sohl-Dickstein et al., 2015), have achieved state-of-the-art results in many generative domains including image (Dhariwal & Nichol, 2021), speech (Popov et al., 2021), video (Ho et al., 2022), and music (Liu et al., 2021) synthesis. Nevertheless, one of the significant drawbacks that distinguishes diffusion-based approaches from other generative models like Variational Autoencoders (Kingma & Welling, 2014) is the lack of an implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations.
To mitigate this issue, several works seek meaningful relations in the approximated starting noise used for generations. This method, known as an inversion technique, was introduced by Song et al. (2021) with Denoising Diffusion Implicit Models (DDIM), and led to the proliferation of works (Garibi et al., 2024; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Samuel et al., ; Hong et al., 2024; Parmar et al., 2023). The core idea is to use the noise predicted by the Diffusion Model and add it to the image instead of subtracting it. Repeating this process effectively traces the backward diffusion trajectory, approximating the noise that could have generated the image. However, due to approximation errors and biases introduced by the trained model, discrepancies arise between the original noise and its reconstruction – latent representation.
@@ 2 Background and Related Work
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal & Nichol (2021) propose to approximate this equation by assuming the local linearity between directions $(x_{t-1}\rightarrow x_{t})$ and $(x_{t}\rightarrow x_{t+1})$, so that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance (Ho & Salimans, 2021; Mokady et al., 2023) for better prompt adherence. As a result, also noticed by recent works (Garibi et al., 2024; Parmar et al., 2023), latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables: Gaussian noise ($\mathbf{x_{T}}$}, an input to generate an image through a backward diffusion process), image sample ($\mathbf{x_{0}}$}, the outcome of the diffusion model generation process), and latent encoding ($\mathbf{\hat{x}_{T}}$}, the result of the DDIM inversion procedure as introduced in Eq. 2).
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting (Zhang et al., 2023a), interpolation (Dhariwal & Nichol, 2021; Zheng et al., 2024), and edition (Su et al., 2022; Kim et al., 2022a; Hertz et al., 2022; Ceylan et al., 2023; Deja et al., 2023). Several works (Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024; Samuel et al., ; Hong et al., 2024; Miyake et al., 2023; Han et al., 2024; Cho et al., 2024; Dong et al., 2023; Zhang et al., 2023b; Parmar et al., 2023; Tang et al., 2024; Wallace et al., 2023; Pan et al., 2023; Wang et al., 2024; Brack et al., 2024; Lin et al., 2024; Ju et al., 2024) aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) (Ho & Salimans, 2021). Null-text inversion (Mokady et al., 2023) extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques improve inversion for image editing through seeking embeddings (Miyake et al., 2023; Han et al., 2024; Dong et al., 2023) or leveraging DDIM latents (Cho et al., 2024) for guidance. On the other hand, some works leverage additional numerical methods (Samuel et al., ; Pan et al., 2023; Garibi et al., 2024) to minimize inversion error. In particular, Renoise (Garibi et al., 2024) iteratively improves the estimation of the next point along the forward diffusion trajectory by averaging multiple noise predictions, incorporating an additional patch-level regularization term that penalizes correlations between pixel pairs to ensure the editability of the latents. Huberman-Spiegelglas et al. (2024) followed by Brack et al. (2024) propose inversion methods for DDPMs, enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. (2023) propose to additionally regularize final DDIM Inversion outputs for better image editing, Lin et al. (2024) introduce an alternative noise scheduler to improve inversion stability, while Hong et al. (2024) propose an exact inversion procedure for higher-order DPM-Solvers, solving the optimization problem at each step.
@@ 3 Analysis
ADM-32 ADM-64 ADM-256 LDM DiT IF
Diffusion Pixel Pixel Pixel Latent Latent Pixel
Image res. $32\times 32$ $64\times 64$ $\times 256 256×256$$\times 256$ $256\times 256$ $64\times 64$
Latent res. - - - $3\times 64\times 64$ $4\times 32\times 32$ -
Trainset CIFAR-10 ImageNet ImageNet CelebA ImageNet LAION-A
Cond?
Backbone U-Net U-Net U-Net U-Net DiT U-Net
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in pixel and latent spaces. More details on models are provided in Appendix G.
In our experiments, we employ six different diffusion models, which we compare in Table 1. For both generation and inversion processes, we use the DDIM sampler with, unless stated otherwise, $T=100$ steps. We provide more details on the number of diffusion steps in Appendix F.
@@ 3.2 Location of latent encodings space
To answer this question, we first geometrically investigate the location of the latents with respect to the generation trajectory. To that end, we analyze the distance between the following steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the DDIM latent. We present the results of this experiment in Fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$.
@@
Similar observation can be made on the basis of visualizations in Fig. 2, where we can distinguish the coarse shape of the original objects in the latents. In particular, while the pixels associated with the objects have high diversity, the areas related to the background are much smoother. This leads to the hypothesis that the most significant difference between the initial noises and latent encodings is the limited variance in the areas corresponding to the background of the generated images. To validate it, we compare the properties of the latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$ that select smooth surfaces in the image, such as sky, sea, plain backgrounds, or surfaces (see Fig. 12 for examples), according to the neighbor differences as described in Appendix I. In Table 3, we show that the error between the starting noise $\mathbf{x_{T}}$} and the latent $\mathbf{\hat{x}_{T}}$} resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across the models, this trend goes along with a decrease in the standard deviation of the latents’ pixels related to those regions. It suggests that DDIM inversion struggles with reversing the plain image areas, bringing them to mean (0) and reducing their diversity.
@@ 3.3 Origin of the divergence
where $\mathcal{E}_{t}^{S}$ is the true model prediction at step $t$, and $\mathcal{E}_{t}^{I}$ is the inversion’s approximation using the previous step’s output. Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the inversion process. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To measure the error solely for the analyzed step $t$, we start the inversion procedure from the exact latent from step (t1)(t-1) (cache from the sampling path). We split the latent pixels into plain and non-plain areas according to the masks calculated for clean images. More details on this setup can be found in Appendix J.
@@ 4 Consequences of the divergence and how to mitigate them?
  • Recent studies (Deja et al., 2022; Liu et al., 2025; Li & Chen, 2024; Fesl et al., 2025) have shown that final steps of the backward diffusion do not contribute additional generative information, instead functioning as a data-agnostic denoising process. Therefore, exact inversion of those steps is less important from the perspective of accurate reconstruction.
The proposed inversion step is therefore defined as follows (see Appendix K for pseudocode):
@@
Before moving to practical applications, we first evaluate the effectiveness of this approach, with $N=10000$ images generated with $T=50$ steps using DiT and IF models. For such generations, we first noise them with a forward diffusion to the intermediate step $t^{\prime}$, followed by the DDIM inversion for $T-t^{\prime}$ steps, ending up with an approximation of the initial noise. In Table 4, we show that by replacing just $4\%$ of the inversion steps, we can completely remove correlations in the latents, up to the level of random Gaussian noise. To further evaluate this effect, in Table 5, we measure the size of the different parts of images (plain vs non-plain) after saving them with the PNG lossless compression. Compression is most effective in the parts related to plain images generated from the DDIM latents, which incline low diversity of their values. At the same time, replacing only 4% of inversion steps significantly reduces this issue. While our simple fix appears to effectively decorrelate the inversion latents, in the following sections, we showcase the consequences of the divergence between noise and latents in several practical use cases.
@@ 4.1 Interpolation quality
We start with the task of image interpolation, where the goal, for two given images, is to generate a sequence of semantically meaningful intermediary frames. Numerous methods (Dhariwal & Nichol, 2021; Song et al., 2021; Samuel et al., 2023; Zheng et al., 2024; Zhang et al., 2024b; Bodin et al., 2025) Diffusion Models with DDIM inversion technique, to calculate latents, interpolate them, and reconstruct the target image. Song et al. (2021) propose to use the spherical linear interpolation (SLERP, Shoemake (1985)), that, for two objects $x$ and $y$, with a coefficient $\lambda\in[0;1]$, is defined as $z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y$, where $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
@@
In our experiment, we compare the quality of interpolations in the noise and latent spaces. To this end, we sample N=20N=20k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into their latents. Next, we randomly assign pairs, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In Fig. 5, we show that, by calculating FID-10k, generations starting from interpolations between random noises (in orange) preserve consistent quality along the entire path – meaning that all interpolated images fall into the real images manifold. In contrast, this property collapses for the setting with DDIM latents as inputs (in purple). In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall (Kynkäänniemi et al., 2019), especially when getting closer to the middle of the interpolation path. Nevertheless, as presented in Fig. 5, using our fix for the first 4% of steps mitigates this issue, enabling higher-quality interpolations with better diversity of the intermediate points. In Section P.2, we present qualitatively that the proposed fix leads to more diverse interpolations, especially in the image background.
@@ 4.2 Diversity and quality of image edition
Apart from image interpolations, text-to-image diffusion models with DDIM inversion are often used for text-based edition, where a source image is first inverted and then reconstructed with a different target prompt (Hertz et al., 2022; Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024). However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting point might reduce the diversity and quality of the edited samples. To evaluate this, we use DiT (Peebles & Xie, 2023) and IF (StabilityAI, 2023) as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used for edition. Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise $\mathbf{x_{T}}$} and invert them back into the latents $\mathbf{\hat{x}_{T}}$}. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. We compare the edits with ground truth targets $I_{T}$ generated from the original noise $\mathbf{x_{T}}$} with $P_{T}$. In Fig. 6, we present the drawback of leveraging latents as starting points for the denoising, where the structures for $I_{S}$ images’ backgrounds limit editing performance in $\hat{I}_{T}$ target images.
@@
Text CLIP-T ($P_{S}$) $\downarrow$ $0.465$ $0.480$ $0.273$ $0.649$
alignment CLIP-T ($P_{T}$) $\uparrow$ $0.681$ $0.662$ $0.353$ $0.614$
Table 6: Diversity of editions (generations from noise $\mathbf{x_{T}}$} and latents $\mathbf{\hat{x}_{T}}$}, conditioned on target prompt) in relation to source images $I_{S}$ and their alignment with source and target prompts. The arrows ($\uparrow$$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
In Table 6, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance (Fu et al., 2023), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and cosine similarity of DINO features (Darcet et al., 2024) to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of Table 6, we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP (Radford et al., 2021) encoder.
Figure 7: Replacing first DDIM inversion steps with forward diffusion improves editions’ pixel diversity (a) and prompt-alignment (b). For IF and DiT models, diversity of generations can be improved by leveraging the forward diffusion process in the first inversion steps, and denoising resulting latents with a different prompt. Additionally, we observe a boost in generations’ alignment to the target prompts, as indicated by the larger CLIP-T value for $P_{T}$ and the smaller one for $P_{S}$. We present details of the experiment in Appendix L.
We further evaluate how replacing the first inversion steps with the forward diffusion affects the diversity and text alignment of generated outputs. As shown in Fig. 7a, swapping the first $4\%$ of DDIM inversion steps with forward diffusion improves the diversity of images generated from latents almost to the level of the samples from original noise. At the same time, as presented in Fig. 7b, replacing the first inversion steps leads to a significant decrease in generations’ alignment to the source prompt ($P_{S}$) and an increase in similarity to the target one ($P_{T}$). This can also be observed in visualizations (Fig. 6). Moreover, because of the random noise added as the initial inversion steps, as presented in Appendix M, our approach enables stochastic image editing, producing diverse manipulations of the same image.

### v5

@@ 1 Introduction
Diffusion-based probabilistic models (DMs), (Sohl-Dickstein et al., 2015), have achieved state-of-the-art results in many generative domains including image (Dhariwal and Nichol, 2021), speech (Popov et al., 2021), video (Ho et al., 2022), and music (Liu et al., 2021) synthesis. Nevertheless, one of the significant drawbacks that distinguishes diffusion-based approaches from other generative models like Variational Autoencoders (Kingma and Welling, 2014) is the lack of an implicit latent space that encodes the images into low-dimensional, interpretable, or editable representations.
To mitigate this issue, several works seek meaningful relations in the approximated starting noise used for generations. This method, known as an inversion technique, was introduced by Song et al. (2021) with Denoising Diffusion Implicit Models (DDIM), and led to the proliferation of works (Garibi et al., 2024; Mokady et al., 2023; Huberman-Spiegelglas et al., 2024; Samuel et al., 2025; Hong et al., 2024; Parmar et al., 2023). The core idea is to use the noise predicted by the Diffusion Model and add it to the image instead of subtracting it. Repeating this process effectively traces the backward diffusion trajectory, approximating the noise that could have generated the image. However, due to approximation errors and biases introduced by the trained model, discrepancies arise between the original noise and its reconstruction – latent representation.
@@ 2 Background and Related Work
However, due to circular dependency on $\epsilon^{(t)}_{\theta}(x_{t},c)$, Dhariwal and Nichol (2021) propose to approximate this equation by assuming the local linearity between directions $(x_{t-1}\rightarrow x_{t})$ and $(x_{t}\rightarrow x_{t+1})$, so that the model’s prediction in $t$-th inversion step can be approximated using $x_{t-1}$ as an input, i.e.,
@@
While such approximation is often sufficient to obtain good reconstructions of images, it introduces the error dependent on the difference $(x_{t}-x_{t-1})$, which can be detrimental for models that sample images with a few diffusion steps or use the classifier-free guidance (Ho and Salimans, 2021; Mokady et al., 2023) for better prompt adherence. As a result, also noticed by recent works (Garibi et al., 2024; Parmar et al., 2023), latents resulting from DDIM inversion do not follow the definition of Gaussian noise because of the existing correlations. In this work, we empirically study this phenomenon and explain its origin. We discuss the relations between the following three variables: Gaussian noise ($\mathbf{x_{T}}$}, an input to generate an image through a backward diffusion process), image sample ($\mathbf{x_{0}}$}, the outcome of the diffusion model generation process), and latent encoding ($\mathbf{\hat{x}_{T}}$}, the result of the DDIM inversion procedure as introduced in Eq. 2).
@@
The DDIM inversion, despite the noise approximation errors, forms the foundation for many applications, including inpainting (Zhang et al., 2023a), interpolation (Dhariwal and Nichol, 2021; Zheng et al., 2024), and edition (Su et al., 2022; Kim et al., 2022a; Hertz et al., 2022; Ceylan et al., 2023; Deja et al., 2023). Several works (Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024; Samuel et al., 2025; Hong et al., 2024; Miyake et al., 2023; Han et al., 2024; Cho et al., 2024; Dong et al., 2023; Zhang et al., 2023b; Parmar et al., 2023; Tang et al., 2024; Wallace et al., 2023; Pan et al., 2023; Wang et al., 2024; Brack et al., 2024; Lin et al., 2024; Ju et al., 2024) aim to reverse the denoising process in text-to-image models, where prompt embeddings can strongly affect the final latent representation through Classifier-free-guidance (CFG) (Ho and Salimans, 2021). Null-text inversion (Mokady et al., 2023) extends the DDIM inversion with additional null-embedding optimization, reducing the image reconstruction error. Other techniques improve inversion for image editing through seeking embeddings (Miyake et al., 2023; Han et al., 2024; Dong et al., 2023) or leveraging DDIM latents (Cho et al., 2024) for guidance. On the other hand, some works leverage additional numerical methods (Samuel et al., 2025; Pan et al., 2023; Garibi et al., 2024) to minimize inversion error. In particular, Renoise (Garibi et al., 2024) iteratively improves the estimation of the next point along the diffusion trajectory by averaging multiple noise predictions, incorporating an additional patch-level regularization term that penalizes correlations between pixel pairs to ensure the editability of the latents. Huberman-Spiegelglas et al. (2024) followed by Brack et al. (2024) propose inversion methods for DDPMs, enabling the creation of various image edition results via inversion. Finally, to reduce the discrepancy between DDIM latents and Gaussian noises, Parmar et al. (2023) propose to additionally regularize final DDIM Inversion outputs for better image editing, Lin et al. (2024) introduce an alternative noise scheduler to improve inversion stability, while Hong et al. (2024) propose an exact inversion procedure for higher-order DPM-Solvers, solving the optimization problem at each step.
@@ 3 Analysis
Model Diffusion Resolution Training Cond? Arch
Space Image Latent Dataset
ADM-32 Pixel 32x32 - CIFAR-10 U-Net
ADM-64 Pixel 64x64 - ImageNet U-Net
ADM-256 Pixel 256x256 - ImageNet U-Net
LDM Latent 256x256 3x64x64 CelebA U-Net
DiT Latent 256x256 4x32x32 ImageNet DiT
IF Pixel 64x64 - LAION-A U-Net
SDXL Latent 1024 4x128x128 - U-Net
Table 1: Overview of diffusion models used for our experiments. We study both unconditioned and conditioned models, operating in pixel and latent spaces. More details on models are provided in Appendix J.
In our experiments, we employ six different diffusion models, which we compare in Table 1. For both generation and inversion processes, we use the DDIM sampler with, unless stated otherwise, $T=100$ steps. We provide more details on the number of diffusion steps in Section N.1.
@@ 3.2 Location of latent encodings space
To answer this question, we first geometrically investigate the location of the latents with respect to the generation trajectory. To that end, we analyze the distance between the following steps $\{x_{t}\}_{t=T\dots 1}$ of the backward diffusion process and intermediate points on the linear interpolation path between the noise and the DDIM latent. We present the results of this experiment in Fig. 3, where each pixel, with coordinates $(t,\lambda)$, is colored according to the $l_{2}$ distance between the intermediate trajectory step $x_{t}$ and the corresponding interpolation step. This distance can be expressed as $\lVert(1-\lambda){\mathbf{x_{T}}}+\lambda{\mathbf{\hat{x}_{T}}}-x_{t}\rVert_{2}$. For better clarity we normalize the distances column-wise.
@@
Similar observation can be made on the basis of visualizations in Fig. 2, where we can distinguish the coarse shape of the original objects in the latents. In particular, while the pixels associated with the objects have high diversity, the areas related to the background are much smoother. This leads to the hypothesis that the most significant difference between the initial noises and latent encodings is the limited variance in the areas corresponding to the background of the generated images. To validate it, we compare the properties of the latents between plain and non-plain areas in the image. We determine binary masks $\mathcal{M}_{p}$ by calculating the absolute difference between neighboring pixels. Pixels where this local variation falls below a fixed threshold ($\tau=0.025$) across all channels are classified as plain regions, effectively isolating low-texture areas (see Appendix H for more details). This procedure results in selection of areas, such as sky, sea, plain backgrounds, or surfaces (see Fig. 12 for examples). In Table 3, we show that the error between the starting noise $\mathbf{x_{T}}$} and the latent $\mathbf{\hat{x}_{T}}$} resulting from the DDIM inversion is higher for pixels corresponding to the plain areas. Across the models, this trend goes along with a decrease in the standard deviation of the latents’ pixels related to those regions. It suggests that DDIM inversion struggles with reversing the plain image areas, bringing them to mean (0) and reducing their diversity. Additionally, in Appendix P, we present that correlations and reduced latent diversity for plain image regions can be similarly observed within Flow matching (Lipman et al., 2023) models.
@@ 3.3 Origin of the divergence
where $\mathcal{E}_{t}^{S}$ is the true model prediction at step $t$, and $\mathcal{E}_{t}^{I}$ is the inversion’s approximation using the previous step’s output. Based on the observations from the previous section, we propose to investigate how the inversion approximation error $\xi(t)$ differs for pixels associated with plain and non-plain image areas throughout the inversion process. To that end, we average the approximation errors for $4000$ images for each of the $T=50$ diffusion timesteps. To measure the error solely for the analyzed step $t$, we start the inversion procedure from the exact latent from step (t1)(t-1) (cache from the sampling path). We split the latent pixels into plain and non-plain areas according to the masks calculated for clean images. More details on this setup can be found in Appendix I.
@@ 4 Consequences of the divergence and how to mitigate them?
  • Recent studies (Deja et al., 2022; Liu et al., 2025; Li and Chen, 2024; Fesl et al., 2025) have shown that final steps of the backward diffusion do not contribute additional generative information, instead functioning as a data-agnostic denoising process. Therefore, exact inversion of those steps is less important from the perspective of accurate reconstruction.
The proposed inversion step is therefore defined as follows (see Appendix F for pseudocode):
@@
Before moving to practical applications, we first evaluate the effectiveness of this approach, with $N=10000$ images generated with $T=50$ steps using DiT and IF models. For such generations, we first noise them with a forward diffusion to the intermediate step $t^{\prime}$, followed by the DDIM inversion for $T-t^{\prime}$ steps, ending up with an approximation of the initial noise. In Table 4, we show that by replacing just $4\%$ of the inversion steps, we can completely remove correlations in the latents, up to the level of random Gaussian noise. This replacement percentage allows us to navigate the trade-off between reconstruction fidelity and latent manipulability. We observe that this trade-off is highly favorable: replacing the first few steps ($t^{\prime}\leq 4\%$) restores the Gaussian properties required for diverse editing, while maintaining a reconstruction error that remains within the perceptual noise floor (see Section N.2 for detailed analysis). To further evaluate this effect, in Table 5, we measure the size of the different parts of images (plain vs non-plain) after saving them with the PNG lossless compression. Compression is most effective in the parts related to plain images generated from the DDIM latents, which incline low diversity of their values. At the same time, replacing only 4% of inversion steps significantly reduces this issue. While our simple fix appears to effectively decorrelate the inversion latents, in the following sections, we showcase the consequences of the divergence between noise and latents in several practical use cases.
@@ 4.1 Interpolation quality
We start with the task of image interpolation, where the goal, for two given images, is to generate a sequence of semantically meaningful intermediary frames. Numerous methods (Dhariwal and Nichol, 2021; Song et al., 2021; Samuel et al., 2023; Zheng et al., 2024; Zhang et al., 2024b; Bodin et al., 2025) Diffusion Models with DDIM inversion technique, to calculate latents, interpolate them, and reconstruct the target image. Song et al. (2021) propose to use the spherical linear interpolation (SLERP, Shoemake (1985)), that, for two objects $x$ and $y$, with a coefficient $\lambda\in[0;1]$, is defined as $z^{(\lambda)}=\frac{\sin{(1-\lambda)\theta}}{\sin{\theta}}x+\frac{\sin{\lambda\theta}}{\sin{\theta}}y$, where $\theta=\arccos\left((x\cdot y)/(\lVert x\rVert\lVert y\rVert)\right)$.
@@
In our experiment, we compare the quality of interpolations in the noise and latent spaces. To this end, we sample N=20N=20k noises, use DDIM with $T=50$ diffusion steps to generate images, and invert those images back into their latents. Next, we randomly assign pairs, which we interpolate with SLERP for $\lambda\in\{0,\frac{1}{6},\frac{1}{3},\frac{1}{2},\frac{2}{3},\frac{5}{6},1\}$ and denoise. In Fig. 5, we show that, by calculating FID-10k, generations starting from interpolations between random noises (in orange) preserve consistent quality along the entire path – meaning that all interpolated images fall into the real images manifold. In contrast, this property collapses for the setting with DDIM latents as inputs (in purple). In the following rows, we show that worse interpolation results for latents stem from the decline in the variety of produced generations, as indicated by a lower recall (Kynkäänniemi et al., 2019), especially when getting closer to the middle of the interpolation path. Nevertheless, as presented in Fig. 5, using our fix for the first 4% of steps mitigates this issue, enabling higher-quality interpolations with better diversity of the intermediate points. In Section O.1, we present qualitatively that the proposed fix leads to more diverse interpolations, especially in the image background.
@@ 4.2 Diversity and quality of image edition
Apart from image interpolations, text-to-image diffusion models with DDIM inversion are often used for text-based edition, where a source image is first inverted and then reconstructed with a different target prompt (Hertz et al., 2022; Mokady et al., 2023; Garibi et al., 2024; Huberman-Spiegelglas et al., 2024). However, knowing that DDIM latents are less diverse in plain areas, we hypothesize that using them as a starting point might reduce the diversity and quality of the edited samples. To evaluate this, we use DiT (Peebles and Xie, 2023) and IF (StabilityAI, 2023) as conditional DMs with $T=50$ diffusion steps. For each model, we construct two sets of $1280$ randomly selected (1) source prompts $P_{S}$, used during the generation and inversion, and (2) target prompts $P_{T}$, used for edition. Using source prompts $P_{S}$, we generate images $I_{S}$ from Gaussian noise $\mathbf{x_{T}}$} and invert them back into the latents $\mathbf{\hat{x}_{T}}$}. Next, we regenerate images $\hat{I}_{T}$ from the latents, changing the conditioning to the target prompts $P_{T}$. We compare the edits with ground truth targets $I_{T}$ generated from the original noise $\mathbf{x_{T}}$} with $P_{T}$. In Fig. 6, we present the drawback of leveraging latents as starting points for the denoising, where the structures for $I_{S}$ images’ backgrounds limit editing performance in $\hat{I}_{T}$ target images.
@@
Text alignment CLIP-T ($P_{S}$) $\downarrow$ $0.465$ $0.480$ $0.273$ $0.353$
CLIP-T ($P_{T}$) $\uparrow$ $0.681$ $0.662$ $0.649$ $0.614$
Directional $\uparrow$ $0.570$ $0.541$ $0.776$ $0.676$
Table 6: Diversity of editions (generations from noise $\mathbf{x_{T}}$} and latents $\mathbf{\hat{x}_{T}}$}, conditioned on target prompt) in relation to source images $I_{S}$ and their alignment with source, target, and directional prompts. The arrows ($\uparrow$$\downarrow$) indicate greater generation diversity and higher text alignment to the target prompt.
In Table 6, we quantitatively measure this effect. First, we calculate the diversity of target generations ($I_{T}$$\hat{I}_{T}$) against source images ($I_{S}$). We use DreamSim distance (Fu et al., 2023), LPIPS (Zhang et al., 2018), SSIM (Wang et al., 2004), and cosine similarity of DINO features (Darcet et al., 2024) to measure the distance between two sets of generations. The experiment shows that edits resulting from latent encodings $\hat{I}_{T}$ are characterized by higher similarity (SSIM, DINO) and lower diversity (DreamSim, LPIPS) relative to starting images $I_{S}$ than the one resulting from noises $I_{T}$. At the same time, in the bottom rows of Table 6, we show that the correlations occurring in the latent encodings induce lower performance in text-alignment to target prompts $P_{T}$, which we measure by calculating cosine similarity between text embeddings and resulting image embeddings, both obtained with the CLIP (Radford et al., 2021) encoder. Additionally, to better assess image editing quality, we use directional CLIP similarity (Gal et al., 2022).
Figure 7: Replacing first DDIM inversion steps with forward diffusion improves editions’ pixel diversity (a) and prompt-alignment (b). For IF and DiT models, diversity of generations can be improved by leveraging the forward diffusion process in the first inversion steps, and denoising resulting latents with a different prompt. Additionally, we observe a boost in generations’ alignment to the target prompts, as indicated by the larger Directional CLIP Similarity, larger CLIP-T value for $P_{T}$, and the smaller one for $P_{S}$. We present details of the experiment in Appendix K.
We further evaluate how replacing the first inversion steps with the forward diffusion affects the diversity and text alignment of generated outputs. As shown in Fig. 7a, swapping the first $4\%$ of DDIM inversion steps with forward diffusion improves the diversity of images generated from latents almost to the level of the samples from original noise. At the same time, as presented in Fig. 7b, replacing the first inversion steps leads to a significant decrease in generations’ alignment to the source prompt ($P_{S}$) and an increase in similarity to the target one ($P_{T}$). This can also be observed in visualizations (Fig. 6). Moreover, because of the random noise added as the initial inversion steps, as presented in Section O.3, our approach enables stochastic image editing, producing diverse manipulations of the same image.